Youve observed the following returns on Yamauchi Corporations stock over the past five years: -27.9 percent, 15.6 percent,
Answer #1

Arithmetic average return=Total return/Total time period

=(-27.9+15.6+34.2+3.3+22.3)/5=9.5%

Return (Return-Average Return)^2
-27.9 (-27.9-9.5)^2=1398.76
15.6 (15.6-9.5)^2=37.21
34.2 (34.2-9.5)^2=610.09
3.3 (3.3-9.5)^2=38.44
22.3 (22.3-9.5)^2=163.84
Total=2248.34%

Standard deviation=[Total (Return-Average Return)^2/(Time period-1)]^(1/2)

=[2248.34/(5-1)]^(1/2)

=[2248.34/4]^(1/2)

=23.71%(Approx).

Variance=Standard deviation^2

=0.056209(Approx).